OpenStudy (anonymous):

Find dy/dx in terms of x and y if x^2 + y^2 = sqrt(8). I just don't know where to begin. Do you find the prime of x^2 + y^2 and then plug in dx/dy? I've been looking online and most people just do the math but don't explain what they are doing.

OpenStudy (anonymous):

\[x^2+y^2=\sqrt{8}\] \[2x+2yy'=0\] solve for \(y'\)

OpenStudy (anonymous):

this is a version of "implicit differentiation" you are thinking that \(y=f(x)\) even though you don't solve for \(y\) in terms of \(x\)

OpenStudy (anonymous):

it is like taking the derivative of \[x^2+f^2(x)=\sqrt{8}\] with respect to \(x\) using the chain rule for the second part you get \[2x+2f(x)f'(x)=0\] but it is easier to write \[2x+2yy'=0\]

OpenStudy (anonymous):

btw you can in fact solve for \(y\) in terms of \(x\) in this problem except you get a \(\pm\) when you do it \[x^2+y^2=\sqrt{8}\] \[y^2=\sqrt{8}-x^2\] \[y=\pm\sqrt{x^2-\sqrt{8}}\]

OpenStudy (anonymous):

Alright let me make sure I get this. (So I got the previous answer) which was -2x / (2y) now with this question: x^3y - x - 7y - 8 = 0 what I would do is first find the derivative and my answer would be: 3x^2y-1 then I plug in y' after the y? 3x^2yy'-1 then I sold for y? 3x^2yy' = 1 y' = 1/(3x^2y)

OpenStudy (anonymous):

\[x^3y - x - 7y - 8 = 0\] \[x^3f(x)-x-7f(x)-8=0\] you need the product rule for the first one

OpenStudy (anonymous):

\[x^3y-x-7y-8=0\] \[3x^2y+x^3y'-1-7y'=0\] solve for \(y'\)